OpenStudy (anonymous):

at what points on the graph of y=1/x is the tangent line perpendicular to 4x-y=3?

OpenStudy (anonymous):

so y=-4x-3 and y=1/x, perpendicular when the derivative of y=1/x is equal to -1/-4(ie gradient of line 2) so when it equals 1/4